# Fractal Dimension in Nonlinear Wave Dynamics Governed by a Nonlinear Partial Differential Equation

### v1

@@ Abstract
keywords:
Boiti–Leon–Pempinelli system; Riccati method; Exact analytical solutions; Fractal structures; Geometric complexity; Self-similarity; Box-counting dimension
@@ 1 Introduction:
Here, $u(x,y,t)$ denotes the primary wave field representing the nonlinear dispersive component, while $v(x,y,t)$ is an auxiliary field coupled to $u$. This system was originally introduced by Boiti, Leon, and Pempinelli in 1987 [1], which generalizes the Korteweg–de Vries (KdV) equation into (2+1) dimensions, incorporating arbitrary spatial and temporal functions. In the limit of reduced dimensionality, the system aligns with the classical dispersive long wave equation, demonstrating its fundamental importance in the theory of nonlinear wave propagation. The system’s mathematical structure has made it a central object of study. Garagash [2] analyzed its Hamiltonian form and identified specific soliton solutions. Later, Lü and Zhang [3] applied the extended tanh method to derive closed-form solutions. Fang et al. [4] expanded this line of research by employing an extended mapping method to reveal fractal and localized patterns in the BLP system. Other significant contributions include Wazwaz and Mehanna [5], who investigated traveling wave solutions using exponential and tanh-coth functions, and Aghdaei and Heris [6], who applied the generalized $(G^{\prime}/G)$-expansion technique. Baskonus and Bulut [7] proposed an improved Bernoulli sub-equation method to construct novel solitary waves. Additional developments include the works of El-Shiekh [8], who studied solutions under variable coefficients, and Yokus et al. [9], who utilized the $(1/G^{\prime})$-expansion method for hyperbolic-type structures. The analytical exploration of the BLP system has been broadened through numerous methodologies encompassing integrability analysis, Lie symmetries, and various functional transformations [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
@@ 2 Construction of Exact Solutions for the BLP System:
[…]
@@ 4 Fractal dimension:
Figures 46 display the results of a voxel-based box-counting analysis applied to fractal surface profiles derived from the BLP system. Each figure showcases a sequence of zoomed-in surface views over progressively smaller domains in the $x$$y$ plane, revealing increasing geometric intricacy at finer spatial scales. As resolution improves, the voxel count needed to cover the surface increases in a nonlinear manner, indicating the presence of self-similar and complex structures. Circular markers highlight the data points extracted from box-counting plots, representing the relationship between box size and the number of boxes required to cover the surface. These markers indicate sampling positions used in the dimension estimation. Collectively, the figures offer compelling visual evidence of fractal scaling behavior within the BLP system, reinforcing the interpretation of its solutions as structurally self-similar across multiple magnification levels.
@@
$x$ $y$ Range Dim
[0.04,0.04][-0.04,0.04][ - 0.04 , 0.04 ] 1.6159
[0.004,0.004][-0.004,0.004][ - 0.004 , 0.004 ] 1.4696
[0.0004,0.0004][-0.0004,0.0004][ - 0.0004 , 0.0004 ] 1.3858
$x$ $y$ Range Dim
[0.04,0.04][-0.04,0.04][ - 0.04 , 0.04 ] 1.7943
[0.004,0.004][-0.004,0.004][ - 0.004 , 0.004 ] 1.5833
[0.0004,0.0004][-0.0004,0.0004][ - 0.0004 , 0.0004 ] 1.5110
$x$ $y$ Range Dim
[0.04,0.04][-0.04,0.04][ - 0.04 , 0.04 ] 1.5403
[0.004,0.004][-0.004,0.004][ - 0.004 , 0.004 ] 1.3921
[0.0004,0.0004][-0.0004,0.0004][ - 0.0004 , 0.0004 ] 1.3185
@@ 6 Acknowledgment:
Financial support provided by the University Grants Commission (UGC), India, is gratefully acknowledged by Saugata Dutta (NTA Ref. No. 211610066362211610066362211610066362) for enabling to fulfill this work.

### v2

@@ Abstract
Keywords: Boiti–Leon–Pempinelli system; Riccati method; Exact analytical solutions; Fractal structures; Geometric complexity; Self-similarity; Box-counting dimension
@@ 1 Introduction:
Here, $u(x,y,t)$ denotes the primary wave field representing the nonlinear dispersive component, while $v(x,y,t)$ is an auxiliary field coupled to $u$. This system was originally introduced by Boiti, Leon, and Pempinelli in 1987 [1], which generalizes the Korteweg–de Vries (KdV) equation into (2+1) dimensions, incorporating arbitrary spatial and temporal functions. In the limit of reduced dimensionality, the system aligns with the classical dispersive long wave equation, demonstrating its fundamental importance in the theory of nonlinear wave propagation. The system’s mathematical structure has made it a central object of study. Garagash [2] analyzed its Hamiltonian form and identified specific soliton solutions. Later, Lü and Zhang [3] applied the extended tanh method to derive closed-form solutions. Fang et al. [4] expanded this line of research by employing an extended mapping method to reveal fractal and localized patterns in the BLP system. Other significant contributions include Wazwaz and Mehanna [5], who investigated traveling wave solutions using exponential and tanh-coth functions, and Aghdaei and Heris [6], who applied the generalized $(G^{\prime}/G)$-expansion technique. Baskonus and Bulut [7] proposed an improved Bernoulli sub-equation method to construct novel solitary waves. Additional developments include the works of El-Shiekh [8], who studied solutions under variable coefficients, and Yokus et al. [9], who utilized the $(1/G^{\prime})$-expansion method for hyperbolic-type structures. The analytical exploration of the BLP system has been broadened through numerous methodologies encompassing integrability analysis, Lie symmetries, and various functional transformations [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20].
@@ 2 Construction of Exact Solutions for the BLP System:
where $a_{1},b_{1},c_{1},a_{2},b_{2},c_{2}$, and $q$ are all regarded as functions of $(x,y,t)$. These were identified by substituting into the governing system Eq. (1) and equating coefficients of like powers of $\zeta$ to zero, yielding the following:
@@ 4 Fractal dimension:
Figures 46 display the results of a voxel-based box-counting analysis applied to fractal surface profiles derived from the BLP system. Each figure showcases a sequence of zoomed-in surface views over progressively smaller domains in the xxyy plane, revealing increasing geometric intricacy at finer spatial scales. As resolution improves, the voxel count needed to cover the surface increases in a nonlinear manner, indicating the presence of self-similar and complex structures. Circular markers highlight the data points extracted from box-counting plots, representing the relationship between box size and the number of boxes required to cover the surface. These markers indicate sampling positions used in the dimension estimation. Collectively, the figures offer compelling visual evidence of fractal scaling behavior within the BLP system, reinforcing the interpretation of its solutions as structurally self-similar across multiple magnification levels.
@@
$x$ & $y$ Range Dim
[0.04,0.04][-0.04,0.04] 1.6159
[0.004,0.004][-0.004,0.004] 1.4696
[0.0004,0.0004][-0.0004,0.0004] 1.3858
$x$ & $y$ Range Dim
[0.04,0.04][-0.04,0.04] 1.7943
[0.004,0.004][-0.004,0.004] 1.5833
[0.0004,0.0004][-0.0004,0.0004] 1.5110
$x$ & $y$ Range Dim
[0.04,0.04][-0.04,0.04] 1.5403
[0.004,0.004][-0.004,0.004] 1.3921
[0.0004,0.0004][-0.0004,0.0004] 1.3185
@@ 6 Acknowledgment:
Financial support provided by the University Grants Commission (UGC), India, is gratefully acknowledged by Saugata Dutta (NTA Ref. No. $211610066362$) for enabling to fulfill this work.